OpenStudy (anonymous):

Evaluate the line integral ydx+x^2dy where c=c1+c2. C1 is the path of the straight line segment from the origin (0,0) to the point (4,23) C2 is the path of the parabola y=-x^2+9x+3 from the point (4,23) to the point (7,17)

OpenStudy (tkhunny):

Should you parameterize the paths?

OpenStudy (anonymous):

When you parametrize you get x(t)=4t and y(t)=23t

OpenStudy (anonymous):

so dx=4dt and dy=23dt

OpenStudy (tkhunny):

Insufficient definition. Specify for what values of t. Also, how about the other one?

OpenStudy (anonymous):

What do you mean? and for C2 i get x(t)=4+3t and y(t)=23-6t

OpenStudy (tkhunny):

Insufficient definition. Specify for what values of t.

OpenStudy (anonymous):

C1: t is between 0 and 4 C2: t is between 4 and 7